OpenStudy (anonymous):

anyone can help me describe how distance ,velocity, accelerations, and energies involved in designing a roller coaster ride.

OpenStudy (agreene):

Well, I'd mention: \[\Sigma\vec{F}=m\gamma\vec{\alpha}\] \[\vec{\alpha}=\int\vec{v}dt\] \[\vec{v}=\int\vec{r}dt\] \[\frac{\Sigma \vec{F}}{m\gamma}=\vec{\alpha}\] \[\int\frac{\Sigma \vec{F}}{m\gamma}dt=\vec{v}\] \[\int\int\frac{\Sigma \vec{F}}{m\gamma}dt=\vec{r}\] and how all of those would influence how much force needs to be contained by the costers track and all that jazz.

OpenStudy (anonymous):

can you explain in words please

OpenStudy (agreene):

Net force = mass * gamma * accel area under accel/time graph = velocity area under velocity/time graph = displacement net force divided by mass*gamma = accel area under net force divided by mass*gamma/time graph = velocity area under area under net force divided by mass*gamma/time = displacement Basically, these are the fundamental relations between these things. As you increase acceleration, you're increasing the net force (from the first eq), for example. So if you're making a coaster, if you make it accelerate really fast, you're gonna have to make the track strong enough to withstand the bigger forces.

OpenStudy (anonymous):

thanks